Three non-perpendicular vectors have magnitudes of 166, 77, 218 Newtons and respectively have angle of 7, 17 and 33 degrees with respect to the x-axis. Find the angle of their resultant.
Round to 2 decimal places.



Answer by Edwin McCravy(20054)


     x-components         y-components
      166∙cos(7°)          166∙sin(7°)
      77∙cos(17°)          77∙sin(17°)
     218∙cos(33°)         218∙sin(33°)
--------------------------------------
sums: 421.2283112          161.4742419

magnitude of resultant = √421.2283112²+161.4742419² =
451.1177462 Newtons.

Angle of resultant, since it is in QI 
since both components are positive:

q = tan-1(161.4742419/421.2283112) = 20.97389537°

Edwin